The problem is about using the addition and multiplication principles together. Here is problem #46:
3(x-2)+2=2(x-4) How do I go about doing this problem?

Answer by stanbon(75887)

operate on the parentheses, to get:
3x-6+2=2x-8
Add some like terms to get:
3x-4=2x-8
Get the "x's" together and the numbers together:
3x-2x=-8+4
x=-4
Cheers,
Stan H.